OpenStudy (anonymous):

What is the exact value of the square root of 38 over 2 the square root of 6.

OpenStudy (dumbcow):

\[\frac{\sqrt{38}}{2\sqrt{6}} = \frac{\sqrt{19}\sqrt{2}}{2\sqrt{3} \sqrt{2}} = \frac{\sqrt{19}}{2\sqrt{3}} = \frac{\sqrt{19}\sqrt{3}}{6}=\frac{\sqrt{57}}{6}\]

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

yw